OpenStudy (anonymous):

what is the solution of 3|x + 5| <= 6?

OpenStudy (anonymous):

\[3\left| x+5 \right|\le6\] \[\left| x+5 \right|\le2\] \[ x+5 \le2 \] or \[x+5\ge-2\] \[x \le-3\] or \[x \ge-7\]

OpenStudy (wasiqss):

x<=-3, and x>=-7

OpenStudy (anonymous):